If vectorA+ vectorB= vectorA=vectorB then find angle btwn vectorA and vectorB

Grade:11

1 Answers

Arun
4 years ago
Dear student
 

The angle between A and B is not defined.

Let vector A = a ; vector B = b.

a+b = a-b, 2b = 0(null vector).

Hence b is a null vector and a is any possible vector.

The catch is that the null vector doesn't have a well defined direction, and hence you cannot find the angle between a and b.